pg-5 (similar to) Question Help * The cost of debt Gronseth Drywall Systems, Inc, is in discussions informed the firm that different maturities will carry different coupon rates and sell at different prices. The firm must ch the bonds wil have a S1,000 par value and tat oncosts will tess per bond The company s taxed at 24% Use the ag pro maton cost of financing with the following alternative (Click on the icon located on the top-right corner of the data table below in order to copy its contents into a with its investment bankers regarding the issuance of new bonds. The t banker has choose among several alternatives. In each case to calculate the after tax Time to 15 $200 Theater-tax cost of financing using the approxrmation formulas % Round to two dec mal places) Enter your answer in the answer box and then click Check Answer. Clear All All parts showing
Answer #1

Cost of debt before tax = (Coupon + (Par value-Net proceeds)/n) divided by (Net proceeds+ Par value)/2

Coupon = 6%*1000 = 60

Net proceeds = par value- discount - flotation cost

=1000-200-35

= 765

Cost of debt before tax =Rd= ( 60+ (1000-765)/15) divided by (765+1000)/2

= 75.67 / 882.5

=8.57%

Cost of debt after tax = Rd*(1-Tax)

= 8.57%*(1- 0.24)

= 6.52%